OpenStudy (anonymous):

price of 9-volt batteries increases acoording to function where t is years after jan 1, 1980. during what year will the price reach $5? P(t)=1.1*e^0.047t price of 9-volt batteries increases acoording to function where t is years after jan 1, 1980. during what year will the price reach $5? P(t)=1.1*e^0.047t @Mathematics

OpenStudy (anonymous):

\[1.1*e^0.047t = 5, e ^{0.047t}=4.55, \ln 4.55=0.047t, 0.047t=1.515 , t=32.2, 1980+32.2=2012.2 \] answer=2012